# Knowledge Migration Framework for Smart Contract Vulnerability Detection

### v1

@@ 3.3.   Data-free knowledge distillation
To ensure that the student model effectively acquires knowledge from the teacher model, we employ a knowledge distillation strategy based on prior knowledge for collaborative training, thereby enhancing the training performance of the student model. The majority of traditional knowledge distillation methods rely on the utilisation of true labels for the purpose of distillation. However, in a data-free distillation scenario, the conventional cross-entropy loss function with true labels is not employed as the target. Instead, the feature outputs derived from a specific layer of the teacher model are employed as prior knowledge, thereby guiding the student in its learning process. The distillation loss and classification loss are calculated using Kullback-Leibler (KL) divergence and mean squared error (MSE), respectively, in order to ascertain the discrepancy between the teacher model’s weights $\hat{P}_{teacher}$ and the student model’s output ${P}_{student}$. This enables the student model to emulate the decision-making process of the teacher during training, thereby capturing the intricate knowledge embedded within the teacher network. By employing a data-free distillation strategy, the student model can enhance its accuracy and robustness without the necessity of real data labels. Ultimately, by weighting and combining the distillation loss and classification loss, the total loss of the student network is obtained, thereby balancing the contributions of both to the model training. These are represented by formulas (1-3). In this context, $\hat{P}_{teacher}$ represents the output of the teacher model at the i-th feature position, C is the number of categories in the task output layer, and $\alpha$ is the weight coefficient that balances the distillation loss and classification loss.
$\mathrm{L_{KL}(\hat{P}_{teacher},P_{student})}=\sum_{i=1}^{n}\hat{P}_{teacher}\log(\frac{\hat{P}_{teacher}}{P_{student}}),\hat{P}_{teacher}=\frac{e^{T(i)}}{\sum_{j=1}^{C}e^{T(j)}}$ (1)
$\mathrm{L_{MSC}(\hat{P}_{teacher},P_{student})}=\frac{1}{C}\sum_{i=1}^{C}(\hat{P}_{teacher}-P_{student})^{2}$ (2)
$\mathrm{L_{concat}=\alpha L_{KL}(\hat{P}_{teacher},P_{student})}+(1-\alpha)\mathrm{L_{MSC}(\hat{P}_{teacher},P_{student})}$ (3)
@@ 3.4.   Adaptive Fusion Module
Typically, standard query mechanisms process the query, keys, and values of input features in a uniform manner, generating weight matrices and weighting and summing them through similarity calculations. The Scaled Dot-Product Attention [34] is an illustrative example. The similarity matrix $QK^{T}$ is computed by the dot product of the query Q and key K. This matrix is scaled by a factor $\sqrt{d_{k}}$ and passed through a softmax function to obtain the attention weights for the keys, where $d_{k}$ is the key dimension. The weighted sum of the values is then calculated as $\frac{QK^{T}}{\sqrt{d_{k}}}*V$, where V is the value matrix. However, the standard query mechanism typically involves single-level feature interaction, which renders it less suitable for downstream vector tasks. As previously stated, we have developed a feature enhancement module that integrates feature grouping modelling with attention distribution calculation. This is known as the Multi-Dimensional Query Enhancement Mechanism. The underlying computational process is illustrated in equations (4-6). In the above equation, $\mathrm{Q^{h},K^{h},V^{h},Z^{h}\in R^{B*(N*K)*C*numhead},X^{{}^{\prime}}\in R^{B*(N*K)*C},W^{0}\in R^{c*c}}$, numhead denotes the number of heads, $Z^{h}$ denotes the output feature matrix of each head, and $\mathrm{X}^{\prime}$ denotes the new aggregation result. Firstly, the input feature matrix is mapped to obtain the query $\mathrm{Q^{(h)}}$, key $\mathrm{K}^{(\mathrm{h})}$, and value $\mathrm{V}^{(\mathrm{h})}$ representations along the channel dimension. The $\mathrm{K}^{(\mathrm{h})}$ and $\mathrm{V}^{(\mathrm{h})}$ are grouped based on a specified number of groups. Within each group, an independent group expansion mechanism is applied to map the features back to the multi-head dimensions, ensuring that different attention heads share the fine-grained feature distribution within the group. Secondly, scaled dot-product attention scores are computed for each group, generating an attention weight matrix based on the matching distribution between queries and keys within the group. This matrix is then normalized and combined to obtain the weighted sum, as shown in equation (5), enabling the adaptive fusion of global information and local relationships. Finally, the output features of each attention head are projected through a projection matrix $W^{0}$ to form a new aggregated representation $\mathrm{X}^{\prime}$.
@@ 4.2.   Experimental setup
Evaluation of indicators. Four evaluation criteria have been selected to assess the performance of the model in the experiment. These are accuracy, recall, precision and F1 value, as illustrated in Equations 11-14.

### v2

@@ 3.3.   Data-free knowledge distillation
To ensure that the student model effectively acquires knowledge from the teacher model, we employ a knowledge distillation strategy based on prior knowledge for collaborative training, thereby enhancing the training performance of the student model. The majority of traditional knowledge distillation methods rely on the utilisation of true labels for the purpose of distillation. However, in a data-free distillation scenario, the conventional cross-entropy loss function with true labels is not employed as the target. Instead, the feature outputs derived from a specific layer of the teacher model are employed as prior knowledge, thereby guiding the student in its learning process.
To enhance the diversity and representativeness of training samples, random noise perturbation is introduced, where the random noise variable serves as the input for generating pseudo-samples. The noise variable is initialized as $\mathrm{z}{\sim}\mathrm{N}(\mu,\sigma^{2})$, where $\mathbb{Z}\in\mathbb{R}^{\mathrm{d}}$, d denotes the dimensionality of the noise variance, and ${\sigma}^{\mathrm{2}}$ represents the variance of the noise, controlling the range of the sampling distribution. Specifically, to ensure that the generated pseudo-samples effectively capture the feature distribution of the teacher model, a noise optimization objective function is introduced. The Mean Squared Error (MSE) is employed to optimize the noise variance z, driving the activation features $\alpha^{i}_{False}$ extracted from the teacher model from pseudo-samples to closely approximate the activation features $\alpha^{i}_{True}$ generated by real data. This optimization is achieved through backpropagation, refining z iteratively. The optimization process is formalized in Equations (1-3), where $g_{t}$ denotes the activation function output, encapsulating prior knowledge, $f_{t}$ represents the mapping function of the teacher model from input to output, $\theta_{t}$ refers to the teacher model parameters, $\eta$ specifies the learning rate controlling the update step size, and $\frac{\alpha_{L_{MSE}}}{\alpha_{z}}$ is the gradient of the noise variance. Notably, z is initialized as a noise term sampled from a Gaussian distribution.
$\mathrm{X_{False}=f_{t}(z;\theta_{t}),\alpha_{False}=g_{t}(X_{False};\theta_{t})}$ (1)
$\mathrm{L_{MSE}=\frac{1}{C}\sum_{i=1}^{c}||\alpha^{i}_{False}-\alpha^{i}_{True}||^{2}}$ (2)
$\mathrm{z\leftarrow z-\eta\frac{\alpha_{L_{MSE}}}{\alpha_{z}}}$ (3)
The distillation loss and classification loss are calculated using Kullback-Leibler (KL) divergence and cross-entropy loss functions (CLF), respectively. To determine the difference between the pseudo-label $\hat{\mathrm{P}}_{\mathrm{teacher}}$ of the teacher model and the output $P_{student}$ of the student model on the optimized pseudo-sample. This enables the student model to emulate the decision-making process of the teacher during training, thereby capturing the intricate knowledge embedded within the teacher network. By employing a data-free distillation strategy, the student model can enhance its accuracy and robustness without the necessity of real data labels. Ultimately, by weighting and combining the distillation loss and classification loss, the total loss of the student network is obtained, thereby balancing the contributions of both to the model training. These are represented by formulas (4-6). In this context, ${P}_{teacher}$ represents the output of the teacher model for the i-th feature position, $\hat{P}_{teacher}$ is the pseudo-label generated by the teacher model, C denotes the feature dimension size, and $\alpha$ is the weight coefficient that balances the distillation loss and classification loss.
$\mathrm{\mathrm{L_{KL}(P_{teacher},P_{student})}=\sum_{i=1}^{n}P_{teacher}\log(\frac{\hat{P}_{teacher}}{P_{student}}),P_{student}=\frac{e^{T(i)}}{\sum_{j=1}^{C}e^{T(j)}}}$ (4)
$\mathrm{\mathrm{L_{CLF}(\hat{P}_{teacher},P_{student})}=-\frac{1}{C}\sum_{i=1}^{C}\hat{P}_{teacher}\mathrm{logP_{student}}^{i}}$ (5)
$\mathrm{\mathrm{L_{concat}=\partial L_{KL}(P_{teacher},P_{student})}+(1-\partial)\mathrm{L_{CLF}(\hat{P}_{teacher},P_{student})}}$ (6)
@@ 3.4.   Adaptive Fusion Module
Typically, standard query mechanisms process the query, keys, and values of input features in a uniform manner, generating weight matrices and weighting and summing them through similarity calculations. The Scaled Dot-Product Attention [34] is an illustrative example. The similarity matrix $QK^{T}$ is computed by the dot product of the query Q and key K. This matrix is scaled by a factor $\sqrt{d_{k}}$ and passed through a softmax function to obtain the attention weights for the keys, where $d_{k}$ is the key dimension. The weighted sum of the values is then calculated as $\frac{QK^{T}}{\sqrt{d_{k}}}*V$, where V is the value matrix. However, the standard query mechanism typically involves single-level feature interaction, which renders it less suitable for downstream vector tasks. As previously stated, we have developed a feature enhancement module that integrates feature grouping modelling with attention distribution calculation. This is known as the Multi-Dimensional Query Enhancement Mechanism. The underlying computational process is illustrated in equations (7-9). In the above equation, $\mathrm{Q^{h},K^{h},V^{h},Z^{h}\in R^{B*(N*K)*C*numhead},X^{{}^{\prime}}\in R^{B*(N*K)*C},W^{0}\in R^{c*c}}$, numhead denotes the number of heads, $Z^{h}$ denotes the output feature matrix of each head, and $\mathrm{X}^{\prime}$ denotes the new aggregation result. Firstly, the input feature matrix is mapped to obtain the query $\mathrm{Q^{(h)}}$, key $\mathrm{K}^{(\mathrm{h})}$, and value $\mathrm{V}^{(\mathrm{h})}$ representations along the channel dimension. The $\mathrm{K}^{(\mathrm{h})}$ and $\mathrm{V}^{(\mathrm{h})}$ are grouped based on a specified number of groups. Within each group, an independent group expansion mechanism is applied to map the features back to the multi-head dimensions, ensuring that different attention heads share the fine-grained feature distribution within the group. Secondly, scaled dot-product attention scores are computed for each group, generating an attention weight matrix based on the matching distribution between queries and keys within the group. This matrix is then normalized and combined to obtain the weighted sum, as shown in equation (5), enabling the adaptive fusion of global information and local relationships. Finally, the output features of each attention head are projected through a projection matrix $W^{0}$ to form a new aggregated representation $\mathrm{X}^{\prime}$.
@@ 4.2.   Experimental setup
Evaluation of indicators. Four evaluation criteria have been selected to assess the performance of the model in the experiment. These are accuracy, recall, precision and F1 value, as illustrated in Equations 14-17.
@@ 5.   Conclusion
Statement
This work has been submitted to the IEEE for possible publication. Copyright may be transferred without notice, after which this version may no longer be accessible.